# Symplectic fillings of unit cotangent bundles of spheres and applications

### v2

@@ 2.1. Closed symplectic manifold containing a symplectic filling
Consider the hyperplane bundle $O_{\mathbb{CP}^{d}}(1)\to\mathbb{CP}^{d}$. Its first Chern class is represented by $\omega_{\operatorname{FS}}$, where $\omega_{\operatorname{FS}}$ denotes the normalized Fubini–Study form on $\mathbb{CP}^{d}$, defined so that
@@
There exists a disk bundle $N(\mathbb{CP}^{d})\subset O_{\mathbb{CP}^{d}}(1)$ equipped with a symplectic form $\Omega_{0}$ such that it restricts to $\omega_{\operatorname{FS}}$ on the zero-section and admits a Liouville vector field, defined away from the zero-section, which is tangent to each fiber and points inward along the boundary; in particular, the boundary $\partial N(\mathbb{CP}^{d})$ is concave. Reversing the orientation of the fibers, $\partial N(\mathbb{CP}^{d})$ becomes the Boothby–Wang bundle associated with the symplectic manifold $(\mathbb{CP}^{d},\omega_{\operatorname{FS}})$; in fact, it is contactomorphic to the standard sphere $(S^{2d+1},\xi_{\operatorname{st}})$.
@@ 2.2. The moduli spaces
$c_{1}(TM)([u])=2(dc_{1}(\mathcal{O}_{\mathbb{CP}^{1}}(1))+c_{1}(\mathcal{O}_{\mathbb{CP}^{1}}(2)))(C)=d+2,$
@@
$2c_{1}(TM)([u])+\frac{\dim X}{2}\cdot\chi(\mathbb{CP}^{1})-\mathrm{codim}(\{p_{0}\}\times\mathbb{CP}^{d})=4d+4.$
@@ 2.3. Compactness of the moduli spaces
Since $C\cdot[\Sigma\times\mathbb{CP}^{d}]=1$, it follows that eaxctly one $a_{j}$ equals $1$ while the others vanish. Without loss of generality, assume this occurs for $j=1$:
@@
Combing this with the inequality (2.3) yields
@@ 3.1. Degree of the evaluation map
Given a point $q_{0}\in(\Sigma\times\mathbb{CP}^{d})\setminus(\{p_{0}\}\times\mathbb{CP}^{d})$, there exists a unique holomorphic sphere $u\in\mathcal{M}$ passing through $q_{0}$.
@@
We denote a curve in question by $u=(u_{1},u_{2})$ with complex curves $u_{1}\colon\mathbb{CP}^{1}\to\Sigma$ and $u_{2}\colon\mathbb{CP}^{1}\to\mathbb{CP}^{d}$. Since $[u]=C$, we have $u_{2}]=0$, which implies that $u_{2}$ is constant. As for $u_{1}$, this is a complex line in $\Sigma\cong\mathbb{CP}^{d}$ passing through the distinct two points. This concludes the uniqueness of holomorphic spheres in question. ∎
@@
Take a point $q_{0}\in\Sigma\times\mathbb{CP}^{d}$ as in the preceding lemma. By that lemma, the fiber $\operatorname{ev}^{-1}(q_{0})$ consists of a single element, namely $\operatorname{ev}^{-1}(q_{0})={[u,z]}$. In view of the splitting of $u^{*}TX$ in (2.2), it follows that $q_{0}$ is a regular value of $\operatorname{ev}$. Hence $\deg(\operatorname{ev})=1$. The assertion concerning the restricted evaluation map then follows directly. ∎
@@ 4.1. Symplectic cobordisms
This is the so-called Brieskorn manifold, known to be diffeomorphic to $S^{2}\times S^{3}$ (see e.g. [DK75, Proposition 6.1]). It also carries the canonical contact structure $\xi_{\ell}$, induced by the standard contact structure on the unit $7$-sphere $S^{7}=\{z\in\mathbb{C}^{4}\mid\|z\|=1\}$. The Brieskorn contact manifold $(Y_{1},\xi_{1})$ is contactomorphic to the unit cotangent bundle $(S^{*}TS^{3},\xi_{\operatorname{can}})$ (see for example [KvK16, Lemma 3.1] and [May09, Lemma 5.3]). A Milnor fiber $V_{\ell}$ of the singularity $z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$ gives an exact symplectic filling of $(Y_{\ell},\xi_{\ell})$. According to [AGLV98, Section 2.7], for $\ell$ and $\ell^{\prime}$ with $\ell<\ell^{\prime}$, the singularity $z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$ is adjacent to $z_{1}^{2\ell^{\prime}}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$; hence [Kea14, Lemma 9.9] leads to an exact symplectic embedding of a Milnor fiber $V_{\ell}$ into the completion of $V_{\ell^{\prime}}$. The latter is diffeomorphic to $V_{\ell^{\prime}}\cup([0,\infty)\times Y_{\ell^{\prime}})$, and the image of the embedding is contained in $V_{\ell^{\prime}}\cup([0,c)\times Y_{\ell^{\prime}})$ for some $c>0$ by compactness. Thus, truncating the completion at the hypersurface $\{c\}\times Y_{\ell^{\prime}}$ and removing the interior of the embedded $V_{\ell}$ shows the existece of an exact symplectic cobordism from $(Y_{\ell},\xi_{\ell})$ to $(Y_{\ell^{\prime}},\xi_{\ell^{\prime}})$. The following theorem tells us that this does not hold for the reverse direction in general.

### v3

@@ 2.1. Closed symplectic manifold containing a symplectic filling
Consider the hyperplane bundle $\mathcal{O}_{\mathbb{CP}^{d}}(1)\to\mathbb{CP}^{d}$. Its first Chern class is represented by $\omega_{\operatorname{FS}}$, where $\omega_{\operatorname{FS}}$ denotes the normalized Fubini–Study form on $\mathbb{CP}^{d}$, defined so that
@@
There exists a disk bundle $N(\mathbb{CP}^{d})\subset\mathcal{O}_{\mathbb{CP}^{d}}(1)$ equipped with a symplectic form $\Omega_{0}$ such that it restricts to $\omega_{\operatorname{FS}}$ on the zero-section and admits a Liouville vector field, defined away from the zero-section, which is tangent to each fiber and points inward along the boundary; in particular, the boundary $\partial N(\mathbb{CP}^{d})$ is concave. Reversing the orientation of the fibers, $\partial N(\mathbb{CP}^{d})$ becomes the Boothby–Wang bundle associated with the symplectic manifold $(\mathbb{CP}^{d},\omega_{\operatorname{FS}})$; in fact, it is contactomorphic to the standard sphere $(S^{2d+1},\xi_{\operatorname{st}})$.
@@ 2.2. The moduli spaces
$c_{1}(TX)([u])=dc_{1}(\mathcal{O}_{\mathbb{CP}^{1}}(1))(C)+c_{1}(\mathcal{O}_{\mathbb{CP}^{1}}(2))(C)=d+2,$
@@
$2c_{1}(TX)([u])+\frac{\dim X}{2}\cdot\chi(\mathbb{CP}^{1})-\mathrm{codim}(\{p_{0}\}\times\mathbb{CP}^{d})=4d+4.$
@@ 2.3. Compactness of the moduli spaces
Since $C\cdot[\Sigma\times\mathbb{CP}^{d}]=1$, it follows that exactly one $a_{j}$ equals $1$ while the others vanish. Without loss of generality, assume this occurs for $j=1$:
@@
Combining this with the inequality (2.3) yields
@@ 3.1. Degree of the evaluation map
Given a point $q_{0}\in(\Sigma\setminus\{p_{0}\})\times\mathbb{CP}^{d}$, there exists a unique holomorphic sphere $u\in\mathcal{M}$ passing through $q_{0}$.
@@
We denote a curve in question by $u=(u_{1},u_{2})$ with complex curves $u_{1}\colon\mathbb{CP}^{1}\to\Sigma$ and $u_{2}\colon\mathbb{CP}^{1}\to\mathbb{CP}^{d}$. Since $[u]=C$, we have $u_{2}]=0$, which implies that $u_{2}$ is constant. As for $u_{1}$, this is a complex line in $\Sigma\cong\mathbb{CP}^{d}$ passing through the two distinct points $p_{0}$ and $q_{01}$, where $q_{01}$ denotes the $\Sigma$-component of $q_{0}\in\Sigma\times\mathbb{CP}^{d}$. This concludes the uniqueness of holomorphic spheres in question. ∎
@@
Take a point $q_{0}\in(\Sigma\setminus\{p_{0}\})\times\mathbb{CP}^{d}$ as in the preceding lemma. By that lemma, the fiber $\operatorname{ev}^{-1}(q_{0})$ consists of a single element, namely $\operatorname{ev}^{-1}(q_{0})={[u,z]}$. In view of the splitting of $u^{*}TX$ in (2.2), it follows that $q_{0}$ is a regular value of $\operatorname{ev}$. Hence $\deg(\operatorname{ev})=1$. The assertion concerning the restricted evaluation map then follows directly. ∎
@@ 4.1. Symplectic cobordisms
This is the so-called Brieskorn manifold, known to be diffeomorphic to $S^{2}\times S^{3}$ (see e.g. [DK75, Proposition 6.1]). It also carries the canonical contact structure $\xi_{\ell}$, induced by the standard contact structure on the unit $7$-sphere $S^{7}=\{z\in\mathbb{C}^{4}\mid\|z\|=1\}$. The Brieskorn contact manifold $(Y_{1},\xi_{1})$ is contactomorphic to the unit cotangent bundle $(S^{*}TS^{3},\xi_{\operatorname{can}})$ (see for example [KvK16, Lemma 3.1] and [May09, Lemma 5.3]). A Milnor fiber $V_{\ell}$ of the singularity $z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$ gives an exact symplectic filling of $(Y_{\ell},\xi_{\ell})$. According to [AGLV98, Section 2.7], for $\ell$ and $\ell^{\prime}$ with $\ell<\ell^{\prime}$, the singularity $z_{1}^{2\ell}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$ is adjacent to $z_{1}^{2\ell^{\prime}}+z_{2}^{2}+z_{3}^{2}+z_{4}^{2}$; hence [Kea14, Lemma 9.9] leads to an exact symplectic embedding of a Milnor fiber $V_{\ell}$ into the completion of $V_{\ell^{\prime}}$. The latter is diffeomorphic to $V_{\ell^{\prime}}\cup([0,\infty)\times Y_{\ell^{\prime}})$, and the image of the embedding is contained in $V_{\ell^{\prime}}\cup([0,c)\times Y_{\ell^{\prime}})$ for some $c>0$ by compactness. Thus, truncating the completion at the hypersurface $\{c\}\times Y_{\ell^{\prime}}$ and removing the interior of the embedded $V_{\ell}$ shows the existence of an exact symplectic cobordism from $(Y_{\ell},\xi_{\ell})$ to $(Y_{\ell^{\prime}},\xi_{\ell^{\prime}})$. The following theorem tells us that this does not hold for the reverse direction in general.